# Generative Visual Instruction Tuning

### v1

@@ 4.3 Main Result
We compare our model, GenLLaVA, against state-of-the-art models on benchmarks for advanced knowledge, general understanding, visual generation, and editing.
As shown in Table 1, GenLLaVA consistently outperforms previous models (GILL, AnyGPT, and MGIE) in most tasks. It achieves higher scores on MathVista (24.9 vs. GILL’s 18.6, AnyGPT’s 24.4, MGIE’s 15.5, Unified-IO 2’s 30.7) and MMMU (29.7 vs. GILL’s 28.8, AnyGPT’s 30.6, MGIE’s 25.6, Unified-IO 2’s 35). For general understanding, GenLLaVA excels in SEED-B (63.5 vs. Unified-IO 2’s 61.8, GILL’s 52.5, AnyGPT’s 44.5, MGIE’s 28.8) and MMB (65 vs. Unified-IO 2’s 71.5, GILL’s 38.2, AnyGPT’s 36, MGIE’s 6.6).
In image editing on the EVR dataset, GenLLaVA scores 64.7, behind MGIE’s 71.5 but ahead of Unified-IO 2 (50.2), GILL (30.4), and AnyGPT (40.3). For visual generation, GenLLaVA has a FID score of 14.3 on CC3M, close to Unified-IO 2 (13.4) and matching AnyGPT. On the CLIP Similarity metric for MS-COCO, GenLLaVA scores 0.71, close to Unified-IO 2’s 0.72 and ahead of GILL (0.67), AnyGPT (0.65), and MGIE (0.66).
While GenLLaVA does not surpass Unified-IO 2 in all categories, it shows strong, balanced performance across various tasks, highlighting its versatility and robustness as a generalist vision-language model with less training time.
@@
We evaluate various models’ performance across diverse visual question-answering datasets, including VQAv2, GQA, VisWiz, TextVQA, and ScienceQA. Our results show that Unified-IO 2 and GenLLaVA consistently perform well across most datasets. Specifically, Unified-IO 2 achieves the highest scores on the ScienceQA (86.2%) and TextVQA (67%), while GenLLaVA demonstrates strong performance on VQAv2 (79.3%) and a competitive score on GQA (62.9%). In contrast, GILL and MGIE exhibit generally lower performance across all datasets, with MGIE notably struggling on VisWiz (20.1%) and TextVQA (26.3%). AnyGPT shows moderate effectiveness, with its best performance on ScienceQA (61%). We used VLMEvalKit [6] to get the results for these datasets, which perform a generation-based evaluation using an LLM judge222 We used GPT-4 (0409) as the judge., the results can be seen in Fig. 3.
@@ 4.5 Ablations
We start with the original instruction tuning dataset from LlaVA-1.5, basically reproducing the original results using a one-stage training recipe. We add the textttLLaVA-Pretrain and generation head to our model; we notice that adding generation capabilities significantly affects the visual understanding capabilities in the model where all the metrics between 2.3% (MM-Vet) to 7.9% (MathVista) points in performance, to compensate for this loss in performance we modify the ration of image generation to image understanding data in our dataset from approximately $\sim$50%-50to $\sim$70%-30%, by adding more image instruction data from LVIS-INSTRUCT4V, LRV-Instruction and other chart understanding datasets. This results in a model with significant generation capabilities that maintain its image-understanding capabilities. We finally add image-generation capabilities using our selected subset of the IPr2IPr dataset. This reduces the image understanding capabilities, but we consider this a small enough change that balances the three tasks while maintaining commendable performance. The results can be seen in Fig. 4a.
@@
The quality of the vision encoder can also have big effects on the final LMM performance. We start by using a CLIP/B, which has the lowest performance, then we compare against a stronger visual encoder and create versions of GenLLaVA, which is trained with CLIP [31] and SigLIP [49], respectively. We can see in Fig. 4b that the SigLIP encoder can achieve generally better performance CLIP encoder. This shows that SigLIP is a better vision encoder for LMM development.
@@
We experiment with varying the number of visual generation tokens, $N$. We noticed that we need more visual generation tokens than GILL ($N=4$) and MGIE ($N=8$) to achieve the best performance. We find that $N=16$ is the best choice for our model. We hypothesize that this is because our model has to balance image generation and editing in the same head and thus needs more visual tokens to capture the complexity of the tasks. The results can be seen in Fig. 3c.
4.6 Comparision with SOTA.
When compared with state-of-the-art models, GenLLaVA maintains similar performance to models of the LlaVA family when evaluated using the average of the scores on the MathVista, MMMU, MMVet, SEED-B, and MMB datasets but lags behind bigger and more specialized models. However, some of these models lack the generative capabilities present in GenLLaVA.
When compared with models of similar size and setup ($\sim 7$b parameters), our model surpasses the original LlaVAv1[25] model by  5% points (37.5% vs. 43.2%), but lags behind LlaVA-1.5 by $\sim$2points (45.3% vs. 43.2%). It is surpassed by the LlaVA-Next [24] family of models by $\sim$7points; by Idefics2 by $\sim$12points (55.7% vs. 43.2%) and by the newly released MiniCPM-Llama3 [43] model by $\sim$17points (60.6% vs. 43.2%). Compared with the absolute state-of-the-art open-source models, GenLLaVA lags behind Yi-VL [45] by $\sim$6points; Emu2 [35] by $\sim$2points, LlaVA-Next [24] (34b) by $\sim$15points; and InternVL 1.5 by $\sim$20points. Compared with the absolute state-of-the-art closed models, GenLLaVA lags behind GPT-4o by $\sim$25points, GPT-4V [1] by $\sim$24points; and the Gemini family [40] of models by  16% points.
@@ 5 Conclusion
In this paper, we have introduced the Generative Large Language, and Visual Assistant (GenLLaVA), a new approach that enables Large Multimodal Models (LMMs) to simultaneously excel in image understanding, generation, and editing without significant loss in individual task performance. We balanced multimodal capabilities in a single model by curating a comprehensive multimodal instruction dataset and developing an innovative single-phase training methodology. Our work sets a new standard for building visual assistants with extra capabilities, and we hope our open-source contributions, including datasets, codebase, and model checkpoints, will serve as valuable resources for the research community, driving further advancements in the field of multimodal AI.

### v2

@@ 4.4 Main Result
We evaluate GenLLaVA against similar models across various tasks, covering advanced knowledge, general understanding, and visual generation tasks. As shown in Table 1, GenLLaVA demonstrates consistent improvements, surpassing models like GILL, AnyGPT, and MGIE in most cases.
For advanced knowledge tasks such as MathVista and MMMU, GenLLaVA performs exceptionally well, achieving the top scores compared to other models. In general understanding, GenLLaVA excels, particularly on SEED-B and MMB, significantly outperforming its competitors. Although MGIE specializes in image editing and maintains a slight edge in this area, GenLLaVA holds its own and outperforms other general models on the EVR dataset.
In visual generation tasks, GenLLaVA shows competitive performance, with results that closely match those of Unified-IO 2 on CC3M and CLIP Similarity while consistently maintaining an edge over other models. These results demonstrate GenLLaVA ’s versatility and ability to achieve superior outcomes across multiple modalities.
GenLLaVA surpasses Unified-IO 2 in visual understanding performance. It shows strong and balanced performance across visual generation tasks, highlighting its versatility and robustness as a generalist vision-language model with significantly less training time.
Figure 5: Comparisons of VQA capabilities among GenLLaVA, Unified-IO 2, MGIE, and GILL. One can observe that GenLLaVA is able to describe the image in detail and respond to commonly asked questions, even addressing the unusual aspects within an image. Hallucinations made by the models are highlighted in red.
@@
We evaluate various models’ performance across diverse visual question-answering datasets, including VQAv2, GQA, VizWiz, TextVQA, and ScienceQA. Our results show that Unified-IO 2 and GenLLaVA consistently perform well across most datasets. Specifically, Unified-IO 2 achieves the highest scores on the ScienceQA (86.2%) and TextVQA (67%), while GenLLaVA demonstrates strong performance on VQAv2 (79.3%) and a competitive score on GQA (62.9%). In contrast, GILL and MGIE exhibit generally lower performance across all datasets, with MGIE notably struggling on VizWiz (20.1%) and TextVQA (26.3%). AnyGPT shows moderate effectiveness, with its best performance on ScienceQA (61%). We used VLMEvalKit from Duan et al. (2024) to get the results for these datasets, which perform a generation-based evaluation using the LLM-as-a-judge protocol. [^2] We used GPT-4 (0409) as the judge. The results can be seen in Fig. 4 and Fig. 5.
@@ 4.6 Ablations
We start with the original instruction tuning dataset from LlaVA-1.5, basically reproducing the original results using a one-stage training recipe. We add the LLaVA-Pretrain and generation head to our model, and we notice that adding generation capabilities significantly affects the visual understanding capabilities, with all metrics degrading between 2.3% (MM-Vet) and 7.9% (MathVista). To compensate for this loss in performance, we modify the ratio of image generation to image understanding data in our dataset from $\sim$50%-50to $\sim$70%-30%, by adding more image instruction data from LVIS-INSTRUCT4V, LRV-Instruction and other chart understanding datasets. This results in a model with significant generation capabilities that maintain its image-understanding capabilities. We finally add image-generation capabilities using our selected subset of the IPr2Pr dataset. This reduces the image understanding capabilities, but we consider this a small enough change that balances the three tasks while maintaining commendable performance. Finally, task tokens are added to the resulting instruction set to condition the model on the desired task. The results can be seen in Table 2(a).
Table 2: Ablation experiments on different datasets. We present evaluation results across various ablation types.
(a) Data Ablation. We study the effect of progressively adding data to the model, starting with visual understanding-only data and then incorporating a mix of both understanding and generation tasks.
Model Variation Advanced Knowledge General Understanding Editing Generation
MathVista MMMU MM-Vet SEED-B MMB EVR CC3M COCO
LLaVA-Finetune 24.7 28.7 30.1 54.7 65.6 - - -
Generation 16.8 27.5 27.8 52.1 59.8 30.2 13.9 0.73
Extra Knowledge 28.2 31.8 32.4 59.7 64.1 28.5 14.0 0.72
IPr2Pr-200K 24.9 29.7 33.1 63.5 65.0 64.7 14.3 0.71
Task Tokens 30.5 37.1 35.8 64.5 66.8 66.9 12.5 0.73
(b) Vision Encoder Ablation. We study the performance of using different vision encoders across several visual understanding and generation benchmarks. We do not condition on the task tokens for this experiment.
Model Variation Advanced Knowledge General Understanding Editing Generation
MathVista MMMU MM-Vet SEED-B MMB EVR CC3M COCO
CLIP/B-224px 23.6 28.6 29.1 53.4 60.3 62.1 15.0 0.68
CLIP/L-336px 24.6 29.2 32.4 59.7 64.1 64.5 14.4 0.70
SigLIP/L-384px 24.9 29.7 33.1 63.5 65.0 64.7 14.3 0.71
(c) Number of Generation Tokens. We study the performance across several visual understanding and generation benchmarks when varying the number of visual generation tokens (denoted as $N$). We do not condition on the task tokens for this experiment.
Model Variation Advanced Knowledge General Understanding Editing Generation
MathVista MMMU MM-Vet SEED-B MMB EVR CC3M COCO
$N=4$ 30.7 35.0 30.5 58.2 64.7 40.3 17.0 0.62
$N=8$ 28.7 32.8 32.4 61.2 65.3 53.3 15.6 0.67
$N=16$ 24.9 29.7 33.1 63.5 65.0 64.7 14.3 0.71
@@
The quality of the vision encoder can also have big effects on the final LMM performance. We start by using a CLIP/B, which has the lowest performance, then we compare against a stronger visual encoder and create versions of GenLLaVA, which is trained with CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023), respectively. We can see in Table 2(b) that the SigLIP encoder generally achieves better performance than the CLIP encoder. This shows that SigLIP is a better vision encoder for LMM development.
@@
We experiment with varying the number of visual generation tokens, $N$, to determine the optimal number required for balancing image generation and understanding tasks. We noticed that we need more visual generation tokens than GILL ($N=4$) and MGIE ($N=8$) to achieve the best performance. We find that $N=16$ is the best choice for our model. We hypothesize that this is because our model has to balance image generation and editing in the same head and thus needs more visual tokens to capture the complexity of the tasks. The results can be seen in Table 2(c).
4.7 Comparison with SOTA.
When compared with state-of-the-art models, GenLLaVA maintains similar performance to models of the LLaVA family when evaluated using the average of the scores on the MathVista, MMMU, MMVet, SEED-B, and MMB datasets. However, it lags behind larger and more specialized models. However, some of these models lack the generative capabilities present in GenLLaVA.
When compared with models of similar size and setup ($\sim 7$b parameters), our model surpasses the original LlaVAv1(Liu et al., 2023b) model by  9% points (37.5% vs. 46.9%), and LlaVA-1.5 (Liu et al., 2024a) by $\sim$1points (45.3% vs. 46.9%). It is surpassed by the LlaVA-Next (Liu et al., 2024b) family of models by $\sim$4points; by Idefics2 (Laurençon et al., 2024) by $\sim$8points (55.7% vs. 46.9%) and by the newly released MiniCPM-Llama3 (Xu et al., 2024) model by $\sim$13points (60.6% vs. 46.9%). Compared with the absolute state-of-the-art open-source models, GenLLaVA lags behind Yi-VL (Young et al., 2024) by $\sim$2points; and surpasses Emu2 (Sun et al., 2024a) by $\sim$1points. It lags behind LlaVA-Next (Liu et al., 2024b) (34b) by $\sim$12points, and InternVL 1.5 (Chen et al., 2024b) by $\sim$15points. Compared with the absolute state-of-the-art closed models, GenLLaVA lags behind GPT-4o by $\sim$26points, GPT-4V (Achiam et al., 2023) by $\sim$20points; and the Gemini family (Gemini Team et al., 2023) of models by  13% points.
@@ 5 Conclusion
In this paper, we have introduced the Generative Large Language and Visual Assistant (GenLLaVA), a comprehensive framework for enabling Large Multimodal Models (LMMs) to excel simultaneously in image understanding, generation, and editing, while maintaining competitive performance. By balancing multimodal capabilities within a single model through the curation of a diverse multimodal instruction dataset and the development of an innovative single-phase training methodology, GenLLaVA sets a new benchmark in the development of multimodal systems. Our results show that unifying generation and understanding under a single framework is possible without compromising their strengths. Our work sets a new standard for building visual assistants with extra capabilities, and we hope our open-source contributions, including datasets, codebase, and model checkpoints, will serve as valuable resources for the research community, driving further advancements in the field of multimodal AI. GenLLaVA’s capabilities can be extended to video understanding, audio-visual tasks, and more advanced real-time multimodal interactions.
@@ Acknowledgements
Appendix A Appendix
A.1 Implementation details
We adopt the hyperparameters in Table A1 for all our experiments.
Table A1: Training Hyperparameters
Hyperparameter Value
Batch Size 128
Max Gradient Norm 1.0
Weight Decay 0.1
Learning Rate 2e-5
Optimizer AdamW
Scheduler Warmup & Cosine Decay
Warmup Ratio 0.03
A.2 Evaluation Details
During the evaluation with VLMEvalKit (Duan et al., 2024), we used the prompt templates listed in Fig A1
Figure A1: Prompt templates. (a) Short VQA includes VQAv2, VizWiz, GQA, and TextVQA and ScienceQA. (b) Long VQA includes MMB, SEED-B, and MM-Vet. (c) Advanced knowledge includes MathVista and MMMU. (d) Generation includes CC3M and MS-COCO. (e) Editing includes EVR. <IMAGE> is the image representation, <QUESTION> denotes each specific question, <ANSWER> is the generated answer, <DESCRIPTION> is an image description for generation or editing, and <GENERATED IMAGE> is the output of the generation head..
A.3 Additional Qualitative Examples
In Tables A2 and A3, we present a qualitative comparative analysis of VQA results between our model, GenLLaVA, and other state-of-the-art models: GPT-4 (Achiam et al., 2023), LLaVA (Liu et al., 2024a), GPT-4o (OpenAI, 2024), and InternVL-1.5 (Chen et al., 2024b). Our model, while smaller than state-of-the-art models, is still able to give detailed and precise responses to given questions, avoiding the introduction of hallucinations, unlike LLaVA-1.5.
Visual input example, Extreme Ironing:
Source: https://www.barnorama.com/wp-content/uploads/2016/12/03-Confusing-Pictures.jpg
User What is unusual about this image?
GenLLaVA (Ours) The unusual aspect of this image is that a man is ironing clothes on the back of a yellow SUV in the middle of a city street. This is not a common sight, as ironing is typically done indoors, in a private setting. The presence of a taxi cab in the background further emphasizes the unexpected nature of the scene, as it juxtaposes the mundane task of ironing with the hustle and bustle of urban life.
User What is unusual about this image?
LLaVA (Liu et al., 2024a) The unusual aspect of this image is that a man is ironing clothes while standing on the back of a moving car. This is not a typical scene, as ironing clothes is usually done indoors, in a stationary position, and not while the vehicle is in motion. The man’s actions are both unconventional and potentially dangerous, as ironing clothes on a moving car can be hazardous and may cause accidents or injuries.
User What is unusual about this image?
GPT-4V (Achiam et al., 2023) The unusual aspect of this image is the man ironing on a portable ironing board mounted to the back of a yellow taxi in the middle of a bustling city street. This activity is not typically associated with such a setting, making it an extraordinary and humorous sight. This represents an extreme form of “extreme ironing,” a performance sport where people iron clothes in unexpected and often difficult locations.
User What is unusual about this image?
GPT-4o (OpenAI, 2024) The image shows a man ironing clothes on an ironing board that is attached to the back of a moving vehicle, which appears to be a yellow taxi. This is highly unusual because ironing is typically done indoors, on stable surfaces, and definitely not while attached to a moving car. This extreme scenario might be an example of “extreme ironing,” a humorous and unconventional sport where people iron clothes in extreme or unusual situations. The image itself blends an ordinary chore with a risky and absurd setting, making it quite striking and humorous.
User What is unusual about this image?
InternVL-1.5 (Chen et al., 2024b) This image depicts an unusual scene in which a person is ironing clothes on an ironing board attached to the back of a taxi while it is parked on a city street. Several elements make this situation peculiar: (1) Location: Ironing is typically done indoors at home or in a laundry room where there are proper facilities for such tasks. (2) Practicality: It’s impractical to set up an ironing board and use an iron outdoors, especially in a busy urban environment where distractions and interruptions could occur frequently.
The juxtaposition of a mundane domestic activity with the bustling cityscape creates a humorous and unexpected scenario.
Table A2: Visual question answering example comparison of GenLLaVA to LLaVA-1.5-7B(Liu et al., 2024a), GPT-4V (Achiam et al., 2023), GPT-4o (OpenAI, 2024), and InternVL-1.5 (Chen et al., 2024b). This table follows the format from Bubeck et al. (2023).
Visual input example, Chicken Nugget Map:
User Can you explain this meme in detail?
GenLLaVA (Ours) The meme is a play on the phrase "beauty is in the eye of the beholder," which means that what one person finds beautiful may not be the same for another. In this case, the person creating the meme finds beauty in the shape of the Earth formed by pieces of fried chicken arranged on a baking tray. It’s a humorous take on appreciating simple things and finding artistic or aesthetic value in everyday objects or food items.
User Can you explain this meme in detail?
LLaVA (Liu et al., 2024a) The meme in the image is a creative and humorous take on food, with a focus on chicken nuggets, as the center of the universe. The meme begins with the phrase “Sometimes I just look at pictures of the Earth from space and I marvel at how beautiful it all is…” The main image features a tray of cooked chicken nuggets, with the chicken pieces positioned to resemble the continents and islands on a world map. The meme playfully suggests that the chicken nuggets represent the Earth, and the various locations depicted in the photo are actually chicken nugget versions of different places. The punchline of the meme appears in smaller text at the bottom, reading “I mean, it’s not the real Earth, but how beautiful it is all is.” This lighthearted and entertaining meme combines food and humor, using the image of chicken nuggets to create a funny and imaginative representation of the world.
User Can you explain this meme?
GPT-4V (Achiam et al., 2023) This meme cleverly uses chicken nuggets shaped like continents to make a humorous comparison. The caption suggests that the person admires pictures of Earth from space for their beauty, but the image shows chicken nuggets on a baking tray arranged to look like a world map. This juxtaposition creates a playful contrast between the expected awe-inspiring images of our planet and the mundane, yet amusing arrangement of food. The humor lies in the unexpected substitution and the notion that even chicken nuggets can evoke a sense of wonder if seen from the right perspective.
User Can you explain this meme in detail?
GPT-4o (OpenAI, 2024) This meme humorously compares the shapes of fried chicken nuggets to a map of the world, as though the person is marveling at the beauty of the Earth from space. The chicken nuggets are arranged in a way that resembles the continents, and the caption plays off the idea of someone gazing at these shapes with awe, as if they’re looking at Earth, rather than just pieces of food.
The humor comes from the exaggerated comparison of something mundane (fried chicken nuggets) to something grand and majestic (the Earth viewed from space). It’s a playful and lighthearted take on how our brains often find patterns in random objects that resemble familiar things.
User Can you explain this meme in detail?
InternVL-1.5 (Chen et al., 2024b) This meme humorously contrasts the profound and contemplative act of admiring the beauty of Earth from space with a more mundane, everyday activity. The top part of the image includes text that says: "Sometimes I just look at pictures of the earth from space and I marvel at how beautiful it all is." Following this statement, instead of showing an image of Earth from space, there is a picture of several pieces of fried chicken arranged on a baking sheet.
The joke lies in the unexpected and humorous twist, where the "beautiful sight" referenced in the text turns out to be something as ordinary and relatable as fried chicken. This creates a light-hearted contrast between grandeur (Earth viewed from space) and simplicity (fried chicken), playing on the idea that both can evoke similar feelings of wonder or satisfaction.
Table A3: Visual question answering example comparison of GenLLaVA to LLaVA-1.5-7B(Liu et al., 2024a), GPT-4V (Achiam et al., 2023), GPT-4o (OpenAI, 2024), and InternVL-1.5 (Chen et al., 2024b). This table follows the format from Bubeck et al. (2023).